# Can a $\xi R\phi^{2}$ non-minimal coupling contribute to the non-relativistic gravitational potential?

### v1

@@ 4 The final result and discussion
Using the above, we obtain from 19 the leading contribution of the non-minimal coupling parameter $\xi$ to the two body gravitational potential,
$V(G^{2},\Lambda,\xi,r)=-\frac{32G^{2}\Lambda\xi m_{1}m_{2}}{\pi r}\left[1-\frac{m_{1}^{2}+m_{2}^{2}}{m^{2}_{1}m^{2}_{2}}\frac{\mu l\hbar}{cr^{2}}\right],$ (20)
where the factor $\mu l\hbar/c$ ensures dimensional consistency, with $l$ being a length scale. In the absence of a cosmological constant, there can be no contribution from $\xi$ to the potential, at least to this order of the perturbation theory. Note that while the right hand side of the momentum space expression 19 is perfectly dimensionally consistent, the Fourier transform of the logarithm necessitates the insertion of such factor. How do we fix $l$? For example, we may set $l=R_{\rm TA,max}$, i.e. the maximum turn around radius defined below 1, as one of the natural length scales in this context, beyond which this result cannot be trusted. However, this choice is not unique and anything concrete beyond this remains elusive to us so far. Note that such ambiguity is absent for the $\xi=0$ part of the potential, due to the appearance of different powers of $G$ in different terms, fixing such term uniquely [3, 4]. (Here we do not have any term first order in $G$).
Note that the second term in the potential diverges if either of the masses is vanishing. If we understand it correctly, we do not expect such behaviour from a non-relativistic quantity. This problem originates from the absence of the scalar rest mass term in the scalar-graviton non-minimal vertex, 6. Does this imply that the $\ln\vec{q}^{2}$ term of 19 should not be related to a physical gravitational potential, or, we simply accept 20 as it is? Once again, any clear reasoning in the favour of either of this remains elusive to us so far.
@@
Further note that the potential of 20, even though corresponds to pure short scale physics, contains two fundamental length scales (apart from $R_{\rm TA,max}$) — the shortest or Planck length corresponding to $G$, and the longest or Hubble horizon scale corresponding to $\Lambda$. We believe this is a novel feature encoded in a short scale physics computation.
It seems an important task to further investigate the effect of $\xi$, as well as the vertex associated with $\Lambda$ on various perturbative quantum gravity phenomenon. As we have emphasised, this coupling is qualitatively different from the minimal ones. It would thus be interesting to investigate the effect of $\xi$ on self energy, various matter-graviton vertices and RG perspectives. A look into the soft theorems might also be interesting in this context. We wish to come to these issues in our future publications.
@@
Acknowledgements : SB would like to thank Amitabha Lahiri and Parthasarathi Majumdar for useful discussions.

### v2

@@ 4 The final result and discussion
Using the above, we obtain from 21, 10, the leading contribution of the non-minimal coupling parameter $\xi$ to the two body long range gravitational potential,
$V(G^{2},\Lambda,\xi,r)=-\frac{32G^{2}\Lambda\xi m_{1}m_{2}}{\pi r}\left[1-\frac{m_{1}^{2}+m_{2}^{2}}{m^{2}_{1}m^{2}_{2}}\frac{1}{r^{2}}\right]$ (22)
We note the appearance of masses in the denominator of the second term on the right hand side. Each mass has to be understood as the inverse of the corresponding Compton wavelength. Thus the first term dominates the second. Perhaps a two loop computation for the class of diagrams that gave the first term would produce a result comparable to the second. Note also that this second term has originated from the Fourier transform of $\ln\vec{q}^{2}$, and any such logarithm originates from the term like $(q^{2}/\mu^{2})^{\epsilon/2}\cdot\epsilon^{-1}$ coming from 9. Hence it is essentially a quantum contribution, whereas the first is not.
In this regard, we note the $\xi=0$ part of the gravitational potential up to one loop [3, 4],
$V(\Lambda=0,r)|_{\rm minimal}=-\frac{Gm_{1}m_{2}}{r}\left[1-\frac{G(m_{1}+m_{2})}{r}-\frac{127G}{30\pi^{2}r^{2}}\right]$ (23)
In 22, we do not have any term of first order in $G$ at all, due to the absence of tree level contribution. In fact the vanishing tree level long range contribution with $\xi\neq 0$ actually saves $\xi$, for otherwise a contribution like $\xi Gm_{1}m_{2}/r$ would have constrained $\xi$ very severely from simple solar system data, as is the case for the parameter of the Brans-Dicke theory [53]. Second, we also note that one can use the $\Lambda$-3 graviton vertex used here to evaluate the two body potential with $\xi=0$. However, the result is expected to be subleading compared to the $\Lambda=0$ ones at our subhorizon scale.
@@
To the best of our knowledge, the vertex associated with $\Lambda$ has not been considered earlier. Note also that the potential of 22 contains two fundamental length scales — the shortest or Planck length corresponding to $G$, and the longest or Hubble horizon scale corresponding to $\Lambda$. We believe this is a novel feature encoded in physics of length scale much small compared to that of the horizon.
Let us now compare the leading term of 22 with that of the ${\cal O}(G^{2}\xi)$ term with $\Lambda=0$ (${\cal O}(r^{-4})$) discussed below 12. The ratio of the first to the second for $m_{2}\gg m_{1}$ behaves as,
$\sim\Lambda r^{2}\cdot\frac{m_{1}}{m_{2}}\cdot m_{1}r$
Recall that a mass term is understood as its associated inverse Compton wavelength ($mc/h$). Imagine now that $m_{2},m_{1}$ are respectively the sun and the Mercury. Taking $\Lambda\sim 10^{-52}{\rm m}^{-2}$ and $r\sim 10^{10}{\rm m}$ to be the radius of the perihellion, the above ratio becomes $\sim 10^{36}$. It is clear that this ratio will keep increasing with increasing $r$ or $m_{1}$ or both. Taking now $m_{1}$ to be the electron, the ratio is instead found to be $10^{-72}$. Thus for ‘large’ $m_{1}$ values, the first term of 23 would dominate over the ${\cal O}(G^{2}\xi)$ terms. We believe the dominating role of a term associated with $\Lambda$ compared to the $r^{-4}$ term for large mass objects, even in the much subhorizon solar scale physics, is interesting in its own right.
Let us now compare 22 with the standard Newton potential for two massive scalars. The leading parts of 22 and the above has a ratio of ${\cal O}(G\Lambda\xi)\sim\xi(L_{P}/L_{C})^{2}$, where $L_{C}\sim\Lambda^{-1/2}$ is the size of the cosmological event horizon and $L_{P}$ is the Planck length. Thus this ratio can at most be $10^{-10}$ in the context of primordial inflationary universe and much lower for our current universe. Also, the ratio of the leading part of 22 and the most subleading part of 23 reads $\sim\Lambda\xi r^{2}$, which is also small in the subhorizon scale we are interested in. We also note that the dependence on $m_{1}$ and $m_{2}$ for the two potentials, 22, 23, are qualitatively very different. This originates from the fact that 22 contains $\Lambda$, which has mass dimension two. Also as we have mentioned towards the end of 1, perturbative quantum gravity predictions in the flat background with a $\Lambda$ are expected to be subleading compared to that of the $\Lambda=0$ case, at least in our current universe where $\Lambda$ is very tiny. Such computations in the early inflationary universe might yield some even non-perturbative, phenomenologically significant contributions, e.g. [36]. Note that in such background, computation of scattering at large scales is not possible due to vacuum instability. Instead the effective action formalism might be useful. We wish to come back to the issue of $\xi R\phi^{2}$ coupling and gravitational potential in the context of early inflationary universe in our future work.
There are further things which seem to need our attention in the context of the matter-graviton non-minimal coupling. For example, a complete evaluation of the two body potential at various post-Minkowski order would be interesting. An investigation on the effect of $\xi$ on gravitational radiation seems to be an important task. These might lead to important constraint on $\xi$, giving us valuable insight about the matter-gravity interaction. It would also be interesting to investigate the effect of $\xi$ on self energy, various matter-graviton vertex functions, and RG perspectives. We wish to address these issues in our future publications.
@@
Acknowledgements : SB would like to thank Amitabha Lahiri and Parthasarathi Majumdar for useful discussions. The authors are indebted to anonymous referees for careful and critical reading of the manuscript and for making many useful comments and suggestions.